OpenStudy (anonymous):

Using the ratio of perfect squares method, what is square root of 108 rounded to the nearest hundredth?

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (bloomlocke367):

okay, given what @amistre64 said, I think you have to find what squares 108 is between, which is 100 and 121. so you do \(\Large\frac{108-100}{121-100}\)

OpenStudy (bloomlocke367):

I think.. I'm probably wrong, one second.

OpenStudy (anonymous):

\[\sqrt{108}= 10.39*10.39=107.9521= 108\]

OpenStudy (bloomlocke367):

okay. that works too, but it's not using the ratio method.

OpenStudy (anonymous):

oh okay idk how to do that

OpenStudy (bloomlocke367):

@amistre64 Can you come help?

OpenStudy (anonymous):

here ill see if @Here_to_Help15 can help

OpenStudy (anonymous):

its okay guys i didnt pass it ive been having trouble all day but o well maybe my teacher can help me tuesday cuz i have a class with her that day

OpenStudy (bloomlocke367):

@amistre64 Can you please help?

OpenStudy (amistre64):

your approach is valid. 8 out of 21 is the ratio between perfect squares 100+8/21 should be a better approx then 100 or 121

OpenStudy (amistre64):

10+8/21 that is

OpenStudy (amistre64):

i worked on a post a few months back maybe where the poster gave a link to the textbook that explained it ..... cant find the post tho

OpenStudy (amistre64):

do we want the solution to sqrt(108) to the hundreths position? or the perfect square approximation, evaluated at the hundreths position?

OpenStudy (amistre64):

10+8/21 = 10.3809...